HLB Health’s current stock price is $50 and its last dividend (D0) was $3.25. HLB Health’s required rate of return is 16%.  

a. What is the growth rate of dividend E(g)?             

b. What is the firm’s expected stock price in 8 years?  

Please show work for answers.
Answer #1

a.

Stock Price = D0(1+g) / (k - g)

50 = 3.25(1+g) / (0.16 - g)

50 * (0.16 - g) = 3.25(1+g)

8 - 50g = 3.25 + 3.25g

8 - 3.25 = 3.25g + 50g

4.75 = 53.25g

g = 4.75/53.25

g = 0.0892 or  8.92%

b.

D9 = D0(1+g)9

= 3.25(1 + 0.0892)9

D8 = 7.012

Let, Expected Stock Price in 8 years = P8

P8 = D9 / (k - g)

= 7.012 / (0.16 - 0.0892)

= 7.012 / 0.0708

P8 = 99.0395 or 99.04